# M2SVid: End-to-End Inpainting and Refinement for Monocular-to-Stereo Video Conversion

### v1

@@ 5 End-to-End Stereoscopic Video Refinement
We introduce M2SVid, an efficient end-to-end model for generating a high-quality right video from the reprojected right video. While this problem is similar to the standard video inpainting task, it comes with a set of unique properties, as described next.
(1)  In our setup where we render a right camera view, the inpainting holes are always introduced in regions where there is an increase in the depth value, as we move horizontally from left to right in the image. Furthermore, the inpainting regions are guaranteed to correspond to the background, i.e. the far plane. Thus, as opposed to a general-purpose inpainting model such as for object removal, our inpainting model can learn to use context from only one side of the inpainting regions, corresponding to the far plane. (2)  In a general inpainting setup, the model is only required to generate the missing pixels, while keeping the rest of the image unaltered. However in our case, the regions outside the inpainting mask can also contain artifacts introduced by errors in the depth map, or interpolation during forward splatting. These artifacts are especially common in presence of thin structures such as fences, where the estimated depth can have substantial errors. Fortunately, we can leverage the original left video to correct the errors introduced by the reprojection operation.
(3)  In a single image inpainting task it is necessary to hallucinate content in the missing regions due to lack of additional information. A video inpainting model can leverage information from multiple frames. However when the inpainting holes are large, the model is still required to perform substantial hallucination. In contrast, the inpainting regions are expected to be much thinner in our task (i.e., as controlled by the maximum disparity $D_{\text{max}}$ parameter). As a result, the inpainting problem is greatly simplified and the model can try to copy information from other temporal frames to avoid hallucination.
@@ 5.1 Overview
An overview of our method is shown in Fig. 3. We base our model on a strong pre-trained video diffusion model to benefit from learned video priors. In particular, we utilize Stable Video Diffusion (SVD) [3], a latent video diffusion model trained to generate videos from an input image. We customize the SVD architecture for the stereoscopic video generation task as follows. First, we condition each latent generation on the input left video $V_{l}$, reprojected right video $V_{r}^{warp}$, as well as the disocclusion masks $M_{r}^{occ}$. Secondly, only for the inpainted pixels, we extend the spatial attention in SVD to operate over all pixels in all frames, instead of just the pixels in the same spatial location. This gives greater flexibility for the model to copy visible pixels from neighboring frames in order to consistently inpaint. Our model is trained end-to-end on stereoscopic datasets to generate the refined right view in a feed-forward manner.
@@ 5.2 High-Frequency Details Preservation
Conventional video inpainting methods utilize masked video to condition inpainting models, a strategy widely adopted in stereo conversion methods [68, 49, 76], where the inpainting model is conditioned solely on the warped right video $V_{r}^{warp}$ and the disocclusion mask $M_{\text{r}}^{\text{occ}}$. However, as discussed in Section 5 (1), the warped right video is likely to contain artifacts even in non-inpainting regions. Thus we propose to also utilize the original left video $V_{\text{l}}$ as an extra input to the model. In more detail, the standard SVD model takes the noisy latents for a video snippet and the VAE encoding of a conditioning image as input. We modify the SVD architecture to instead take the following inputs as conditioning: the VAE encoding of the left video snippet $E(V_{l})$, the VAE encoding of the warped right snippet $E(V_{r}^{warp})$, and the disocclusion mask $M_{\text{r}}^{occ,resized}$, resized to the same resolution as the VAE encodings. We thus denote the model $\hat{\boldsymbol{\mathbf{v}}}_{\theta}(\boldsymbol{\mathbf{z}}_{t},t,c)$ with conditioning inputs $c$ as
@@
where [.][.][ . ] refers to concatenation. This modification is achieved by modifying the first convolution layer in the U-Net, as is the common practise.
Using the original left video as conditioning allows to model to easily infer high frequency details as well as other information which might have been destroyed during the depth based reprojection stage. As shown in Fig. 6, this improves the quality of the generated right views, with high-frequency details from the left view better preserved.
@@ 5.3 Spatio-Temporal Aggregation for Inpainting
The SVD model is trained to take multiple video frames (i.e. a snippet) as input and jointly denoise them in order to generate a temporally consistent output. Ideally, one would like to compute a full attention over all spatial tokens in all frames to enable maximal information flow. However, this operation would be prohibitively expensive, resulting in a complexity of $H=N^{2}\times h^{2}\times w^{2}$ for each attention layer, where $h$ and $w$ are the size of the latent representations. For this reason, the standard video diffusion models such as SVD and VideoCrafter [5] only factorize the full attention computation into interleaved spatial and temporal attention layers, to propagate information across multiple locations and frames. In the spatial attention layers, only tokens from the same timestamp attend to each other, while in the temporal attention layers, tokens from the same spatial location attend to each other across different timestamps. This drastically reduces the complexity to the order of only $N\times h^{2}\times w^{2}+N^{2}\times h\times w$, which is more palatable.
The factorization of the full attention computation is however likely to reduce the modelling capacity for our task, specially in dynamic scenes with camera motion. In such cases, as discussed in Sec. 5 (2), an occluded pixel in one frame is likely visible at a different spatial location in another frame, simplifying the inpainting problem. In such cases, we would like to perform a full spatio-temporal attention. Fortunately, the number of inpainted tokens constitute only a small fraction of all tokens (usually $<$ 5%). We exploit this property and modify the spatial attention layer in SVD to allow tokens corresponding to the disocclusion mask $M^{occ}_{r}$ to attend to all other tokens, while using the spatial attention only for other tokens. This improves the inpainting capability of our model (Fig. 5) without any significant computational overhead.
@@ 5.4 Efficient Feed-Forward Prediction
The diffusion models for text-to-image/video generation and inpainting commonly utilize multiple denoising steps to generate the final image [20, 62, 60]. This is important due to the ill-posed nature of the problem wherein multiple solutions exists for a single input prompt. Utilizing fewer denoising steps usually results in blurry output.
However, in our task of right view generation, the inpainting regions are generally small. Furthermore, as described in Sec. 5, the neighboring temporal frames as well as the input left frames contain a significant amount of information needed to generate a high-quality, temporally consistent right video. The model therefore does not need to perform significant hallucination in most cases, but rather needs to fetch the relevant information from other frames. The relatively constrained nature of our problem thus motivates us to generate the right view in a direct feed-forward manner, similar to the approach employed by Garcia et al. [16] for depth estimation.
Given the input left video $V_{\text{l}}$, warped right video $V_{\text{r}}^{warp}$, and the disocclusion masks $M_{\text{r}}^{occ}$, we generate the conditioning input $c$ to the model by obtaining the VAE encodings of left and reprojected videos, as shown in Eq. 3. The initial latent is set to the mean noise 0 as in [16]. The conditioning input, together with the initial latent are then passed through the Video U-Net to obtain the latent encoding $\hat{z}$ of the right video. This is then passed through the VAE decoder $D$ to obtain the predicted right view $\hat{V}_{r}$. The inference pipeline can thus be denoted as,
@@
Here $z$ corresponds to the VAE encodings of the ground truth right video $V_{r}$. Note that during the training, keep the VAE decoder frozen and only fine tune the U-Net model.
@@ 5.5 Training Data
In order to train our models using the supervised objective Eq. 5, we require datasets consisting of stereoscopic videos, together with the disparity maps between the stereo pairs. Due to the lack of large-scale stereoscopic datasets, prior works have focused on collecting and processing private datasets from the internet, which usually have not been publicly released. Instead we aim to train using publicly available datasets.
We utilize Ego4D [17] and Stereo4D [25] datasets for our training. Ego4D a large-scale egocentric video dataset containing approximately 263 long videos (80 hours in total) collected using a stereo camera, while the recently released Stereo4D [25] dataset consists of $\sim$ 200K stereoscopic video clips sourced from $\sim$ 7K online videos. The Stereo4D dataset also provides rectified videos and disparity maps. The Ego4D dataset on the other hand contains unrectified videos, without any precomputed disparity maps.
We thus perform the following steps to preprocess Ego4D dataset in order to use it for training. First, we uniformly sample $K=200$ frames from the video and perform dense feature matching using LoFTR [71]. These are used to compute the fundamental matrix [91] using RANSAC [12]. Next, we estimate the rectification transformations for both left and right frames. After rectifying the images, we spatially crop the video to discard boundary regions with missing pixels. We then compute the LoFTR matches again and shift the left and right views horizontally until all disparities between the matched points are positive and the smallest disparity is zero. This ensures that the left and right videos follow a rectified stereo setup. Finally, we utilize an off-the-shelf stereo matching method BiDaVideo [26] to obtain dense disparity maps for all stereo pairs.
Figure 4: Qualitative comparison of our approach with state-of-the-art methods SVG [10] and StereoCrafter [68]. Our approach can effectively preserve the high-frequency information from the input video and generate high-quality right views.
@@ 6 Experimental Results
Quantitative evaluation:  To quantitatively evaluate our approach, we need datasets containing left and right stereo videos, together with the (pseudo) ground-truth disparity maps which are used to generate the warped right views for each method. This allows us to directly compare the generated right views with the ground truth right views, without having to worry about errors introduced by incorrect depth estimation. We rely on the test sets of the Stereo4D [25] and Ego4D [17] datasets, processed as described in Sec 5.5. As metric, we use the standard image quality metrics PSNR, MS-SSIM, and LPIPS [89] to compare the different methods. We compute the metrics independently over each video, and average them over the full dataset. In the appendix, we also report the metrics computed only inside, as well as outside the disoccluded regions. All evaluations are performed on 16-frame videos, sampled at 5 FPS, resized to 512 resolution, and centrally cropped.
Qualitative evaluation:  For qualitative analysis and user studies, we only need monocular videos along with per-frame depth maps. We use videos from the DAVIS dataset [57] as well as videos from free online stock sources [^1] https://www.pexels.com/.We sample 16 frames per video at 8 FPS and resize them to $640\times 1152$ resolution. We compute the depth maps using the recent DepthCrafter [22] method.
@@
SVG [10] - (training free) 50 steps 25.6 0.926 0.217
StereoCrafter [68] private dataset 25 steps 24.9 0.909 0.242
StereoCrafter [68] private dataset 1 step 25.3 0.911 0.262
@@ 6.1 Comparison to State-of-the-Art
Baselines:  We compare our method to the recent state-of-the-art monocular-to-stereo conversion models SVG [10] and StereoCrafter [68]. SVG is a training-free method that utilizes a frozen diffusion model to inpaint regions within the disocclusion mask while not modifying the remaining regions. StereoCrafter, which is an unpublished concurrent work, fine-tunes SVD [3] to inpaint and refine the warped right views in a diffusion-based denoising setup, using the warped right views along with the disocclusion masks as conditioning. We evaluate StereoCrafter using the default 25 denoising steps as well as single denoising step. We exclude the Deep3D [82] from the evaluation as it doesn’t allow control over the baseline between the stereo cameras and is shown to have inferior results [68, 10] compared to SVG and StereoCrafter. In order to ensure a fair comparison, we use the same depth maps for reprojection in all methods. However, we use the official warping implementation provided by the authors for each method.
@@
Qualitative Results:  We perform a qualitative comparison with the SVG and StereoCrafter methods in 4. Since SVG only performs inpainting within the disocclusion mask, it fails to fix the artifacts introduced by errors in warping. Furthermore, it can incorrectly extend the ‘foreground’ object to fill the inpainting hole due to lack of task-specific fine tuning. StereoCrafter with a single denoising step produces blurry inpainting and degrades the quality of the warped areas as well. This is expected since the model was trained using a multi-step denoising objective. When using 25 denoising steps, StereoCrafter inference generally produces sharp results. However it can struggle at times to correctly generate the high-frequency details. This is because the high-frequency information can often get degraded during the warping step. Since StereoCrafter only utilizes warped view and masks as conditioning, it can struggle to recover the details. In contrast, our model M2SVid is conditioned on the input left view as well, allowing it to leverage the full context for inpainting and refinement. Furthermore, since our method is trained in an end-to-end manner with image space losses, it can learn to minimize the loss of high-frequency information introduced by VAE decoder.
Method Denoising steps Average rank $\downarrow$ Latency (s) $\downarrow$
SVG [10] 50 steps 2.88 1270.4
StereoCrafter [68] 25 steps 2.05 12.2
StereoCrafter [68] 1 step 3.46 2.4
M2SVid (Ours) 1 step 1.43 2.1
Table 2: Average human perception study rankings over 21 videos with 13 distinct users. The ranking ranges from 1 to 4, with 1 being the best. Our method significantly outperforms StereoCrafter with an average rank of 1.43. We also show the run-times computed using an A100 GPU on a $512\times 512$ videos with 16 frames.
Quantitative Results:  We quantitatively compare our approach with SVG and StereoCrafter on Stereo4D in Tab. 1. Despite our method M2SVid performing only a single step inference, it significantly outperforms StereoCrafter on all metrics and obtains the best PSNR and LPIPS scores.
@@
User study:  We also perform a user study to evaluate the quality of the generated right views. We use 21 videos containing 16 frames each for the study. The videos are sourced from DAVIS dataset, as well as publicly available sources. For each video, we show the right videos generated by each method (anonymized). Additionally, we also show the input left video, and the disocclusion mask as reference. Each of the 13 participants were asked to rank the quality of the generated videos from 1 (best) to 4 (worst), taking into account factors such as temporal consistency, image quality (i.e. sharpness) and lack of artifacts. We collected 112 rankings per method. The average rankings are summarized in Tab. 2. Our method significantly outperforms all others, obtaining an average ranking of 1.43, compared to 2.05 obtained by StereoCrafter with 25 denoising steps.
Figure 5: Using full-attention at dis-occluded pixels (Sec. 5.3) enables the model to exploit information from visible pixels in other frames to improve inpainting.
@@
Run-time:  According to Tab. 2, on an A100 GPU, our approach M2SVid achieves a 6$\times$ and 635$\times$ speed-up compared to StereoCrafter with 25 steps and SVG respectively.

### v2

@@ 5 End-to-End Stereoscopic Video Refinement
We introduce M2SVid, an efficient end-to-end model for generating a high-quality right video from the warped right video $V_{r}^{warp}$. While this problem is similar to the standard video inpainting task, it comes with a set of unique properties, as described next.
(1)  In our setup for rendering a right camera view, inpainting holes appear in regions where depth increases from left to right. These regions correspond to background, so unlike general-purpose inpainting (e.g., object removal), our model can rely on context from only one side of the hole.
(2)  In a general inpainting setup, the model is only required to generate the missing pixels, while keeping the rest of the image unaltered. In our case, however, regions outside the inpainting mask may also contain artifacts from depth-map errors or interpolation during forward splatting. These artifacts are especially common in presence of thin structures such as fences, where the estimated depth can have substantial errors. Fortunately, we can leverage the original left video to correct the errors introduced by the reprojection.
(3)  In the image inpainting task, the model must hallucinate content in missing regions due to the lack of additional information. A video inpainting model can instead use information from multiple frames, though large holes still require substantial hallucination. In our task, however, the inpainting regions are much thinner (i.e., as controlled by the maximum disparity $D_{\text{max}}$). As a result, the inpainting problem is greatly simplified and the model can copy information from other temporal frames to avoid hallucination.
@@ 5.1 Overview
An overview of our method is shown in Fig. 3. We base our model on a strong pre-trained video diffusion model to benefit from learned video priors. In particular, we utilize Stable Video Diffusion (SVD) [Blattmann2023StableVD], a latent video diffusion model trained to generate videos from an input image. We customize the SVD architecture for the stereoscopic video generation task as follows. First, we condition each latent generation on the input left video $V_{l}$, reprojected right video $V_{r}^{warp}$, as well as the disocclusion masks $M_{r}^{occ}$. Secondly, only for the inpainted pixels, we extend the spatial attention in SVD to operate over all pixels in all frames, instead of just the pixels in the same spatial location. This gives greater flexibility for the model to copy visible pixels from neighboring frames in order to consistently inpaint. Finally, we train our model end-to-end on public stereoscopic datasets to generate the refined right view in a feed-forward manner, without requiring multiple denoising steps.
@@ 5.2 High-Frequency Details Preservation
Conventional video inpainting methods utilize masked video to condition inpainting models, a strategy widely adopted in stereo conversion methods [shi2024stereocrafter, mehl2024stereo, wang2023learning], where the inpainting model is conditioned solely on the warped right video $V_{r}^{warp}$ and the disocclusion mask $M_{\text{r}}^{\text{occ}}$. However, as discussed in Sec. 5 (2), the warped right video is likely to contain artifacts even in non-inpainting regions. Thus we propose to also utilize the original left video $V_{\text{l}}$ as an extra input to the model. In more detail, the standard SVD model takes the noisy latents for a video snippet and the VAE encoding of a conditioning image as input. We modify the SVD architecture to instead take the following inputs as conditioning: the VAE encoding of the left video snippet $E(V_{l})$, the VAE encoding of the warped right snippet $E(V_{r}^{warp})$, and the disocclusion mask $M_{\text{r}}^{occ,resized}$, resized to the same resolution as the VAE encodings. We thus denote the model $\hat{\boldsymbol{\mathbf{v}}}_{\theta}(\boldsymbol{\mathbf{z}}_{t},t,c)$ with conditioning inputs $c$ as
@@
where $[.]$ refers to concatenation. This modification is achieved by modifying the first convolution layer in the U-Net, as is the common practise.
Using the original left video as conditioning allows the model to easily infer high frequency details as well as other information which might have been destroyed during the depth based reprojection stage. As shown in Fig. 5, this improves the quality of the generated right views, with high-frequency details from the left view better preserved.
@@ 5.3 Spatio-Temporal Aggregation for Inpainting
The SVD model takes multiple video frames (i.e. a snippet) as input and jointly denoise them to produce a temporally consistent output. Ideally, one would compute a full attention over all spatial tokens in all frames for maximal information flow. However, this would be prohibitively expensive, resulting in a complexity of $H=N^{2}\times h^{2}\times w^{2}$ for each attention layer, where $h$ and $w$ are the size of the latent representations. For this reason, the standard video diffusion models such as SVD and VideoCrafter [chen2024videocrafter2] factorize the full attention computation into interleaved spatial and temporal attention layers. In the spatial attention layers, only tokens from the same timestamp attend to each other, while in the temporal attention layers, tokens from the same spatial location attend to each other across different timestamps. This drastically reduces the complexity to $N\times h^{2}\times w^{2}+N^{2}\times h\times w$, which is more palatable.
However, factorizing full attention can reduce modeling capacity for our task, especially in dynamic scenes with camera motion. As discussed in Sec. 5(3), an occluded pixel in one frame is often visible at a different spatial location in another, simplifying inpainting problem. In such cases, we would prefer full spatio-temporal attention. Fortunately, the number of inpainted tokens constitute only a small fraction of all tokens (usually << 5%). We exploit this property and modify the spatial attention layer in SVD to allow tokens corresponding to the disocclusion mask $M^{occ}_{r}$ to attend to all other tokens, while using the spatial attention only for other tokens. This improves the inpainting capability of our model (Fig. 6) without significant computational overhead.
Figure 4: Qualitative comparison of our approach with state-of-the-art methods SVG [dai2024svg] and StereoCrafter [shi2024stereocrafter]. Our approach can effectively preserve the high-frequency information from the input video and generate high-quality right views.
@@ 5.4 Efficient Feed-Forward Prediction
The diffusion models for text-to-image/video generation and inpainting commonly utilize multiple denoising steps to generate the final image [ho2020denoising, saharia2022photorealistic, rombach2022high]. This is important due to the ill-posed nature of the problem wherein multiple solutions exists for a single input prompt. Utilizing fewer denoising steps usually results in blurry output. However, in our task of right view generation, the inpainting regions are generally small. Furthermore, as described in Sec. 5, the neighboring temporal frames as well as the input left frames contain a significant amount of information needed to generate a high-quality, temporally consistent right video. The model therefore does not need to perform significant hallucination in most cases, but rather needs to fetch the relevant information from other frames. The relatively constrained nature of our problem thus motivates us to generate the right view in a direct feed-forward manner, rather than the multi-step denoising used in prior work [dai2024svg, shi2024stereocrafter], similar to the approach employed by Garcia et al[garcia2024fine] for depth estimation.
Given the input left video $V_{\text{l}}$, warped right video $V_{\text{r}}^{warp}$, and the disocclusion masks $M_{\text{r}}^{occ}$, we generate the conditioning input $c$ to the model by obtaining the VAE encodings of left and reprojected videos, as shown in Eq. 3. The initial latent is set to the mean noise 0 as in [garcia2024fine] and the timestamp is set to the highest noise $t=T$. As discussed in Sec. 3, this choice of initial latent and timestamp leads the Video U-Net to directly reconstruct a clean latent due to the use of v-parameterization [salimans2022progressive]. Therefore, the conditioning input, together with the initial latent are then passed through the Video U-Net $\hat{\boldsymbol{\mathbf{v}}}_{\theta}$ to obtain the latent encoding $\hat{z}=-\hat{\boldsymbol{\mathbf{v}}}_{\theta}(\textbf{0},T,c)$ of the right video. This is then passed through the VAE decoder $D$ to obtain the predicted right view $\hat{V}_{r}$. The inference pipeline can thus be denoted as,
@@
Here $z$ corresponds to the VAE encodings of the ground truth right video $V_{r}$. Note that during training, we keep the VAE decoder frozen and only fine-tune the U-Net model.
@@ 5.5 Training Data
To train our model using the supervised objective Eq. 5, we require a dataset stereoscopic videos with their corresponding disparity maps. Due to the lack of large-scale stereoscopic datasets, prior works focused on collecting private datasets that usually have not been publicly released. Instead we aim to train using publicly available datasets.
We utilize Ego4D [grauman2022ego4d] and Stereo4D [jin2024stereo4d] datasets for our training. Ego4D a large-scale egocentric video dataset containing approximately 263 long videos (80 hours in total) collected using a stereo camera, while the recently released Stereo4D [jin2024stereo4d] dataset consists of $\sim$ 200K stereoscopic video clips sourced from $\sim$ 7K online videos. The Stereo4D dataset also provides rectified videos and disparity maps. The Ego4D dataset on the other hand contains unrectified videos, without any precomputed disparity maps.
We thus perform the following steps to preprocess Ego4D dataset. First, we uniformly sample frames and perform dense feature matching with LoFTR [sun2021loftr], using these matches to compute the fundamental matrix [Zhang2021] via RANSAC [Fishchler]. Next, we estimate rectification transformations for both views and rectify the videos. We then compute the LoFTR matches again and shift the left and right views horizontally until all disparities between the matched points are positive and the smallest disparity is zero. This ensures that the videos follow a rectified stereo setup. Finally, we use an off-the-shelf stereo matching method BiDaVideo [jing2024match-stereo-videos] to obtain disparity maps for all stereo pairs.
@@ 6 Experimental Results
Quantitative evaluation:  To quantitatively evaluate our approach, we need datasets containing left and right stereo videos, together with the (pseudo) ground truth disparity maps which are used to generate the warped right views for each method. This allows us to directly compare the generated right views with the ground truth right views, without having to worry about errors introduced by incorrect depth estimation. We rely on Stereo4D [jin2024stereo4d] and Ego4D [grauman2022ego4d] test sets and use the standard image quality metrics PSNR, MS-SSIM, and LPIPS [zhang2018perceptual] for evaluation. We compute the metrics independently over each video, and average them over the full dataset. In the appendix, we also report metrics inside and outside disoccluded regions. All evaluations are performed on 16-frame videos, sampled at 5 FPS, resized to 512 resolution, and centrally cropped.
@@
SVG [dai2024svg] - (training free) 50 steps 25.6 0.926 0.217
StereoCrafter [shi2024stereocrafter] private dataset 25 steps 24.9 0.909 0.242
StereoCrafter [shi2024stereocrafter] private dataset 1 step 25.3 0.911 0.262
@@
[…]
steps (rank=1) $\uparrow$
SVG [dai2024svg] 50 steps 2.88 16 / 112 1270.4
StereoCrafter [shi2024stereocrafter] 25 steps 2.05 29 / 112 12.2
StereoCrafter [shi2024stereocrafter] 1 step 3.46 4 / 112 2.4
M2SVid (Ours) 1 step 1.43 76 / 112 2.1
[…]
VR headset comparison StereoCrafter [shi2024stereocrafter] (25 steps) Ours (1 step)
preferred preferred
StereoCrafter vs. Ours 9 / 105 39 / 105
Table 3: VR headset human perception study with 105 comparisons over 21 videos with 5 distinct users. Our method shows a clear advantage over StereoCrafter with 25 denoising steps.
Qualitative evaluation:  For qualitative analysis and user studies, we only need monocular videos along with per-frame depth maps. We use videos from the DAVIS dataset [perazzi2016benchmark] as well as videos from free online stock sources [^2] https://www.pexels.com/. We sample 16 frames per video at 8 FPS and resize them to $640\times 1152$ resolution. We compute the depth maps using the recent DepthCrafter [hu2024depthcrafter] method.
@@ 6.1 Comparison to State-of-the-Art
Baselines:  We compare our method to the recent state-of-the-art monocular-to-stereo conversion models SVG [dai2024svg] and StereoCrafter [shi2024stereocrafter]. SVG is a training-free method that utilizes a frozen diffusion model to inpaint regions within the disocclusion mask while not modifying the remaining regions. StereoCrafter, which is an unpublished concurrent work, fine-tunes SVD 1.1 [Blattmann2023StableVD] to inpaint and refine the warped right views in a diffusion-based denoising setup, using the warped right views along with the disocclusion masks as conditioning. We evaluate StereoCrafter using the default 25 denoising steps as well as single denoising step. We exclude the Deep3D [xie2016deep3d] from the evaluation as it doesn’t allow control over the baseline between the stereo cameras and is shown to have inferior results [shi2024stereocrafter, dai2024svg] compared to SVG and StereoCrafter. In order to ensure a fair comparison, we use the same depth maps for reprojection in all methods. However, we use the official warping implementation provided by the authors for each method.
@@
Qualitative Results:  We perform a qualitative comparison with the SVG and StereoCrafter methods in Fig. 4. Since SVG only performs inpainting within the disocclusion mask, it fails to fix the artifacts introduced by errors in warping. Furthermore, it can incorrectly extend the ‘foreground’ object to fill the inpainting hole due to lack of task-specific fine-tuning. StereoCrafter with a single denoising step produces blurry inpainting and degrades the quality of the warped areas as well. This is expected since the model was trained using a multi-step denoising objective. When using 25 denoising steps, StereoCrafter inference generally produces sharp results. However it can struggle at times to correctly generate the high-frequency details. This is because the high-frequency information can often get degraded during the warping step. Since StereoCrafter only utilizes warped view and masks as conditioning, it can struggle to recover the details. In contrast, our model is conditioned on the input left view as well, allowing it to leverage the full context for inpainting and refinement. Furthermore, since our method is trained in an end-to-end manner with image space losses, it can learn to minimize the loss of high-frequency information introduced by VAE decoder.
Quantitative Results:  We quantitatively compare our approach with SVG and StereoCrafter on Stereo4D in Tab. 1. Despite using only a single-step inference, M2SVid significantly outperforms both baselines on all metrics, except for SVG on MS-SSIM, where SVG benefits in MS-SSIM from preserving non-empty pixels (warped using ground truth disparity), while our method incurs a slight MS-SSIM drop due to VAE compression, but achieves superior visual quality by correcting warping errors, e.g., in thin structures.
@@
User studies:  We also perform two user studies. First, in a desktop user study, participants viewed a random subset of 21 videos from the DAVIS dataset and public sources. For each video, anonymized right views from each method were shown alongside the input left view and disocclusion mask (as reference). Each of the 13 participants were asked to rank the quality of the generated videos from 1 (best) to 4 (worst), taking into account factors such as temporal consistency, image quality (i.e. sharpness) and lack of artifacts. If methods are indistinguishable, the equal ranking was allowed. In total, 112 rankings per method were collected (Tab. 2). Our method significantly outperforms all others, obtaining an average ranking of 1.43, compared to 2.05 obtained by StereoCrafter (25 denoising steps) and 2.88 by SVG. In fact, our method was ranked first $2.6\times$ more often than StereoCrafter (25 steps), and $4.75\times$ than SVG.
To validate that the enhanced visual quality of our method also improves user experience in stereoscopic viewing, we conducted a second user study using a VR headset. Participants were shown anonymized stereo videos generated by our method and StereoCrafter (25 denoising steps) and asked which version they preferred, or if both were equal. We collected 105 comparisons from 5 users (21 each). As shown in Tab. 3, our method was preferred in 39 cases, while StereoCrafter was favored in only 9 (the rest are tied), highlighting the clear advantage of our approach.
@@
Figure 6: Using full-attention at dis-occluded pixels (Ours,  Sec. 5.3) enables the model to exploit information from visible pixels in other frames to improve inpainting (see Appendix).
[…]